Matthew is 3 times as old as jenny. In 7 years, he will be twice as old as she will be then. How old is each now?

Answer by checkley77(12844)

M=3x
J=x
In 7 years:
(3x+7)=2(x+7)
3x=7=2x+14
3x-2x=14-7
x=7 Jenny's age now.
3*7=21 Matthew's age now.
Proof:
3*7+7=2(7+7)
21+7=2*14
28=28